# Context is Key: A Benchmark for Forecasting with Essential Textual Information

### v1

@@ C.2 Extended results on all models
     ARIMA 0.480 $\pm$ 0.006 19.780 $\pm$ 0.386
     ETS 0.522 $\pm$ 0.009 24.259 $\pm$ 0.476
     Exp-Smoothing 0.603 $\pm$ 0.013 25.858 $\pm$ 0.241
C.3 Significant failures per model
We observe that in a few instances in the benchmark, some models tend to obtain significantly worse performance when evaluated with context. In our evaluation, we term all instances where the RCRPS value of a model is greater than $5$, as significant failures of the model on those instances. We found $5$ as a suitable value for analyzing such failures, as it intuitively represents the value a forecast would get if the distance between the forecast and the ground-truth was 5 times bigger than the range of the ground-truth for the task. When we aggregate the RCRPS of instances in the benchmark (such as in Tab. 1), we cap the RCRPS of such significant failures to $5$, to avoid outliers with a much higher RCRPS affecting the aggregate score. In Tab. 4, we show the number of such instances in our evaluation of the benchmark where we found models to have significant failures (out of a total of $355$ evaluated instances). Interestingly, some models such as Direct Prompt with Llama-3.1-405B-Instruct and LLMP with Llama-3-70B and Llama-3-8B are more robust to such significant failures, and do not incur such failures. On the other hand, models such as LLMP with Llama-3-70B-Inst and Llama-3-8B-Inst achieve significant failures in a large number of instances from the benchmark. We postulate that this is because of models misinterpreting context. It is still an open question as to how to increase the robustness of models to prevent or reduce such significant failures. We visualize such significant failures in Sec. C.5.
@@ C.3 Significant failures per model
Direct Prompt (ours)
     Llama-3.1-405B-Inst 0
     Llama-3-70B-Inst 1
     Llama-3-8B-Inst 3
     Mixtral-8x7B-Inst 5
     GPT-4o 5
     GPT-4o-mini 2
@@
     Llama-3-70B-Inst 18
     Llama-3-70B 0
     Llama-3-8B-Inst 12
     Llama-3-8B 0
     Mixtral-8x7B-Inst 1
     Mixtral-8x7B 1

### v2

@@ C.2 Full results partitioned by types of context
    ARIMA 0.480 $\pm$ 0.006 25.582 $\pm$ 0.450 0.565 $\pm$ 0.010 0.200 $\pm$ 0.007 0.332 $\pm$ 0.003 0.386 $\pm$ 0.006 0.440 $\pm$ 0.011
    ETS 0.522 $\pm$ 0.009 31.838 $\pm$ 0.603 0.627 $\pm$ 0.014 0.362 $\pm$ 0.014 0.295 $\pm$ 0.005 0.397 $\pm$ 0.010 0.508 $\pm$ 0.017
    Exp-Smoothing 0.603 $\pm$ 0.013 34.048 $\pm$ 0.333 0.700 $\pm$ 0.020 0.493 $\pm$ 0.016 0.378 $\pm$ 0.006 0.485 $\pm$ 0.016 0.827 $\pm$ 0.060
Sec. C.2 provides the results of all tested models, partitioned by the types of context.
C.3 Full results partitioned by model capabilities
We provide an additional view of the results of all models in Sec. C.3, partitioned by model capabilities.
Table 4: Results of all models on the CiK benchmark. Starting from the left, the first column shows the RCRPS averaged over all tasks. The second column shows the rank of each method w.r.t. other baselines, averaged over all tasks. The remaining columns show the average RCRPS stratified by model capabilities. All averages are weighted according to the scheme described in Sec. 5.1 and accompanied by standard errors. Lower is better and the best averages are in bold.
{tabu}
lccccccccc Average RCRPS Average Rank Common-Sense Retrieval Reasoning
With Context
direct prompt (ours)
    Llama-3.1-405B-Inst 0.159 $\pm$ 0.008 7.440 $\pm$ 0.586 0.140 $\pm$ 0.013 0.109 $\pm$ 0.002 0.191 $\pm$ 0.006 0.133 $\pm$ 0.001 0.167 $\pm$ 0.008 0.316 $\pm$ 0.028 0.376 $\pm$ 0.039
    Llama-3-70B-Inst 0.286 $\pm$ 0.004 14.081 $\pm$ 0.195 0.323 $\pm$ 0.008 0.122 $\pm$ 0.003 0.408 $\pm$ 0.012 0.168 $\pm$ 0.002 0.492 $\pm$ 0.019 0.473 $\pm$ 0.012 0.577 $\pm$ 0.017
    Llama-3-8B-Inst 0.461 $\pm$ 0.008 26.913 $\pm$ 0.559 0.323 $\pm$ 0.010 0.174 $\pm$ 0.003 0.849 $\pm$ 0.021 0.406 $\pm$ 0.014 1.245 $\pm$ 0.039 0.437 $\pm$ 0.004 0.494 $\pm$ 0.004
    Mixtral-8x7B-Inst 0.523 $\pm$ 0.023 31.389 $\pm$ 0.388 0.433 $\pm$ 0.043 0.204 $\pm$ 0.000 0.864 $\pm$ 0.029 0.427 $\pm$ 0.024 1.245 $\pm$ 0.006 0.644 $\pm$ 0.080 0.789 $\pm$ 0.112
    Qwen-2.5-7B-Inst 0.289 $\pm$ 0.003 21.816 $\pm$ 0.724 0.343 $\pm$ 0.005 0.127 $\pm$ 0.002 0.324 $\pm$ 0.008 0.204 $\pm$ 0.005 0.281 $\pm$ 0.014 0.409 $\pm$ 0.002 0.480 $\pm$ 0.002
    Qwen-2.5-1.5B-Inst 0.575 $\pm$ 0.014 30.636 $\pm$ 0.834 0.610 $\pm$ 0.007 0.214 $\pm$ 0.004 0.988 $\pm$ 0.049 0.341 $\pm$ 0.009 1.077 $\pm$ 0.122 0.896 $\pm$ 0.006 1.151 $\pm$ 0.008
    Qwen-2.5-0.5B-Inst 0.464 $\pm$ 0.012 26.261 $\pm$ 0.513 0.268 $\pm$ 0.008 1.029 $\pm$ 0.055 0.744 $\pm$ 0.039 0.244 $\pm$ 0.007 2.043 $\pm$ 0.104 0.330 $\pm$ 0.004 0.392 $\pm$ 0.005
    GPT-4o 0.276 $\pm$ 0.010 8.601 $\pm$ 0.410 0.180 $\pm$ 0.004 0.087 $\pm$ 0.003 0.519 $\pm$ 0.029 0.113 $\pm$ 0.006 0.447 $\pm$ 0.029 0.590 $\pm$ 0.033 0.769 $\pm$ 0.046
    GPT-4o-mini 0.353 $\pm$ 0.022 16.461 $\pm$ 0.462 0.296 $\pm$ 0.043 0.419 $\pm$ 0.014 0.471 $\pm$ 0.012 0.218 $\pm$ 0.005 1.024 $\pm$ 0.033 0.475 $\pm$ 0.080 0.578 $\pm$ 0.112
llmp
    Llama-3-70B-Inst 0.550 $\pm$ 0.013 17.043 $\pm$ 0.535 0.645 $\pm$ 0.018 0.284 $\pm$ 0.015 0.392 $\pm$ 0.014 0.519 $\pm$ 0.026 0.312 $\pm$ 0.019 0.453 $\pm$ 0.020 0.495 $\pm$ 0.028
    Llama-3-70B 0.237 $\pm$ 0.006 11.960 $\pm$ 0.668 0.310 $\pm$ 0.011 0.126 $\pm$ 0.009 0.217 $\pm$ 0.007 0.134 $\pm$ 0.003 0.241 $\pm$ 0.019 0.294 $\pm$ 0.008 0.329 $\pm$ 0.010
    Llama-3-8B-Inst 0.484 $\pm$ 0.010 18.216 $\pm$ 0.443 0.345 $\pm$ 0.002 0.138 $\pm$ 0.004 0.910 $\pm$ 0.030 0.242 $\pm$ 0.008 1.278 $\pm$ 0.069 0.617 $\pm$ 0.022 0.787 $\pm$ 0.030
    Llama-3-8B 0.313 $\pm$ 0.023 18.358 $\pm$ 0.947 0.404 $\pm$ 0.043 0.124 $\pm$ 0.003 0.280 $\pm$ 0.026 0.179 $\pm$ 0.014 0.267 $\pm$ 0.015 0.530 $\pm$ 0.084 0.661 $\pm$ 0.117
    Mixtral-8x7B-Inst 0.264 $\pm$ 0.004 15.210 $\pm$ 0.605 0.344 $\pm$ 0.004 0.127 $\pm$ 0.003 0.224 $\pm$ 0.005 0.179 $\pm$ 0.010 0.173 $\pm$ 0.009 0.348 $\pm$ 0.005 0.405 $\pm$ 0.007
    Mixtral-8x7B 0.262 $\pm$ 0.008 15.890 $\pm$ 0.481 0.348 $\pm$ 0.012 0.146 $\pm$ 0.022 0.230 $\pm$ 0.016 0.153 $\pm$ 0.002 0.230 $\pm$ 0.041 0.354 $\pm$ 0.007 0.414 $\pm$ 0.009
    Qwen-2.5-7B-Inst 1.980 $\pm$ 0.027 42.505 $\pm$ 0.715 1.825 $\pm$ 0.048 1.022 $\pm$ 0.054 2.215 $\pm$ 0.046 2.761 $\pm$ 0.024 1.723 $\pm$ 0.092 2.025 $\pm$ 0.093 1.607 $\pm$ 0.127
    Qwen-2.5-7B 0.904 $\pm$ 0.037 35.488 $\pm$ 0.975 0.690 $\pm$ 0.063 0.794 $\pm$ 0.083 1.558 $\pm$ 0.062 0.879 $\pm$ 0.028 2.328 $\pm$ 0.153 0.878 $\pm$ 0.084 0.881 $\pm$ 0.113
    Qwen-2.5-1.5B-Inst 2.153 $\pm$ 0.027 47.555 $\pm$ 0.835 2.052 $\pm$ 0.046 1.566 $\pm$ 0.033 2.671 $\pm$ 0.038 2.156 $\pm$ 0.035 3.635 $\pm$ 0.053 2.480 $\pm$ 0.085 2.323 $\pm$ 0.113
    Qwen-2.5-1.5B 1.731 $\pm$ 0.036 42.324 $\pm$ 0.515 1.343 $\pm$ 0.061 1.737 $\pm$ 0.074 2.594 $\pm$ 0.042 2.256 $\pm$ 0.042 3.275 $\pm$ 0.132 2.036 $\pm$ 0.083 1.526 $\pm$ 0.114
    Qwen-2.5-0.5B-Inst 1.938 $\pm$ 0.024 47.508 $\pm$ 0.586 1.743 $\pm$ 0.043 1.800 $\pm$ 0.021 2.193 $\pm$ 0.025 2.303 $\pm$ 0.028 3.439 $\pm$ 0.004 1.685 $\pm$ 0.084 1.398 $\pm$ 0.114
    Qwen-2.5-0.5B 1.991 $\pm$ 0.024 45.726 $\pm$ 0.797 1.827 $\pm$ 0.045 0.950 $\pm$ 0.025 1.967 $\pm$ 0.020 2.799 $\pm$ 0.022 1.804 $\pm$ 0.036 1.695 $\pm$ 0.085 1.443 $\pm$ 0.113
Multimodal Models
    UniTime 0.370 $\pm$ 0.001 33.377 $\pm$ 0.151 0.267 $\pm$ 0.002 0.179 $\pm$ 0.001 0.321 $\pm$ 0.001 0.511 $\pm$ 0.003 0.337 $\pm$ 0.001 0.333 $\pm$ 0.001 0.385 $\pm$ 0.001
    Time-LLM (ETTh1) 0.476 $\pm$ 0.001 41.641 $\pm$ 0.142 0.448 $\pm$ 0.002 0.192 $\pm$ 0.000 0.373 $\pm$ 0.000 0.538 $\pm$ 0.003 0.397 $\pm$ 0.001 0.382 $\pm$ 0.001 0.440 $\pm$ 0.001
Without Context
direct prompt (ours)
    Llama-3.1-405B-Inst 0.473 $\pm$ 0.005 32.450 $\pm$ 0.278 0.393 $\pm$ 0.002 0.325 $\pm$ 0.006 0.752 $\pm$ 0.015 0.494 $\pm$ 0.009 0.720 $\pm$ 0.027 0.594 $\pm$ 0.006 0.617 $\pm$ 0.006
    Llama-3-70B-Inst 0.470 $\pm$ 0.008 33.486 $\pm$ 0.348 0.404 $\pm$ 0.002 0.304 $\pm$ 0.007 0.717 $\pm$ 0.015 0.488 $\pm$ 0.022 0.694 $\pm$ 0.024 0.606 $\pm$ 0.007 0.631 $\pm$ 0.007
    Llama-3-8B-Inst 0.601 $\pm$ 0.006 39.657 $\pm$ 0.383 0.391 $\pm$ 0.005 0.322 $\pm$ 0.004 1.123 $\pm$ 0.018 0.642 $\pm$ 0.012 1.446 $\pm$ 0.035 0.581 $\pm$ 0.005 0.617 $\pm$ 0.005
    Mixtral-8x7B-Inst 0.712 $\pm$ 0.021 44.240 $\pm$ 0.357 0.625 $\pm$ 0.043 0.324 $\pm$ 0.006 1.053 $\pm$ 0.007 0.783 $\pm$ 0.005 1.237 $\pm$ 0.007 0.858 $\pm$ 0.080 0.872 $\pm$ 0.112
    Qwen-2.5-7B-Inst 0.420 $\pm$ 0.022 35.132 $\pm$ 0.508 0.447 $\pm$ 0.043 0.259 $\pm$ 0.008 0.505 $\pm$ 0.010 0.374 $\pm$ 0.007 0.411 $\pm$ 0.016 0.692 $\pm$ 0.080 0.792 $\pm$ 0.112
    Qwen-2.5-1.5B-Inst 0.450 $\pm$ 0.007 34.470 $\pm$ 0.592 0.377 $\pm$ 0.006 0.232 $\pm$ 0.004 0.661 $\pm$ 0.018 0.387 $\pm$ 0.013 0.939 $\pm$ 0.039 0.423 $\pm$ 0.007 0.476 $\pm$ 0.008
    Qwen-2.5-0.5B-Inst 0.305 $\pm$ 0.007 25.473 $\pm$ 0.325 0.267 $\pm$ 0.010 0.162 $\pm$ 0.001 0.384 $\pm$ 0.008 0.300 $\pm$ 0.011 0.440 $\pm$ 0.016 0.315 $\pm$ 0.002 0.367 $\pm$ 0.002
    GPT-4o 0.441 $\pm$ 0.008 29.858 $\pm$ 0.359 0.381 $\pm$ 0.002 0.179 $\pm$ 0.002 0.692 $\pm$ 0.028 0.356 $\pm$ 0.007 0.953 $\pm$ 0.067 0.422 $\pm$ 0.003 0.471 $\pm$ 0.002
    GPT-4o-mini 0.423 $\pm$ 0.006 34.039 $\pm$ 0.264 0.360 $\pm$ 0.003 0.214 $\pm$ 0.002 0.649 $\pm$ 0.019 0.392 $\pm$ 0.008 0.771 $\pm$ 0.041 0.461 $\pm$ 0.003 0.511 $\pm$ 0.003
llmp
    Llama-3-70B-Inst 0.378 $\pm$ 0.004 24.665 $\pm$ 0.445 0.369 $\pm$ 0.003 0.150 $\pm$ 0.003 0.513 $\pm$ 0.012 0.292 $\pm$ 0.006 0.668 $\pm$ 0.025 0.384 $\pm$ 0.003 0.440 $\pm$ 0.003
    Llama-3-70B 0.312 $\pm$ 0.006 20.965 $\pm$ 0.435 0.349 $\pm$ 0.011 0.141 $\pm$ 0.002 0.351 $\pm$ 0.008 0.216 $\pm$ 0.004 0.395 $\pm$ 0.017 0.372 $\pm$ 0.004 0.434 $\pm$ 0.005
    Llama-3-8B-Inst 0.503 $\pm$ 0.009 29.628 $\pm$ 0.420 0.385 $\pm$ 0.003 0.159 $\pm$ 0.002 0.914 $\pm$ 0.030 0.430 $\pm$ 0.010 1.271 $\pm$ 0.070 0.424 $\pm$ 0.003 0.467 $\pm$ 0.001
    Llama-3-8B 0.345 $\pm$ 0.003 23.990 $\pm$ 0.356 0.326 $\pm$ 0.004 0.150 $\pm$ 0.002 0.497 $\pm$ 0.009 0.267 $\pm$ 0.005 0.640 $\pm$ 0.017 0.381 $\pm$ 0.002 0.444 $\pm$ 0.002
    Mixtral-8x7B-Inst 0.383 $\pm$ 0.015 23.531 $\pm$ 0.439 0.357 $\pm$ 0.003 0.550 $\pm$ 0.072 0.459 $\pm$ 0.047 0.211 $\pm$ 0.008 1.027 $\pm$ 0.133 0.371 $\pm$ 0.004 0.428 $\pm$ 0.004
    Mixtral-8x7B 0.306 $\pm$ 0.007 21.673 $\pm$ 0.469 0.360 $\pm$ 0.013 0.146 $\pm$ 0.002 0.327 $\pm$ 0.008 0.202 $\pm$ 0.005 0.340 $\pm$ 0.016 0.382 $\pm$ 0.002 0.445 $\pm$ 0.003
    Qwen-2.5-7B-Inst 1.023 $\pm$ 0.026 37.612 $\pm$ 1.177 0.520 $\pm$ 0.044 1.157 $\pm$ 0.022 1.634 $\pm$ 0.017 1.072 $\pm$ 0.035 3.319 $\pm$ 0.037 0.853 $\pm$ 0.081 0.769 $\pm$ 0.112
    Qwen-2.5-7B 0.732 $\pm$ 0.030 35.801 $\pm$ 1.101 0.649 $\pm$ 0.048 0.359 $\pm$ 0.036 0.974 $\pm$ 0.037 0.752 $\pm$ 0.039 1.433 $\pm$ 0.096 0.728 $\pm$ 0.084 0.730 $\pm$ 0.112
    Qwen-2.5-1.5B-Inst 1.515 $\pm$ 0.033 44.209 $\pm$ 1.056 1.314 $\pm$ 0.055 1.159 $\pm$ 0.057 1.802 $\pm$ 0.031 1.653 $\pm$ 0.042 3.383 $\pm$ 0.106 1.108 $\pm$ 0.082 0.848 $\pm$ 0.112
    Qwen-2.5-1.5B 1.069 $\pm$ 0.028 38.510 $\pm$ 1.025 1.005 $\pm$ 0.048 0.287 $\pm$ 0.026 1.339 $\pm$ 0.023 1.264 $\pm$ 0.040 1.798 $\pm$ 0.047 0.771 $\pm$ 0.086 0.720 $\pm$ 0.112
    Qwen-2.5-0.5B-Inst 1.318 $\pm$ 0.037 41.697 $\pm$ 0.715 1.463 $\pm$ 0.064 0.239 $\pm$ 0.031 1.192 $\pm$ 0.019 1.434 $\pm$ 0.047 1.675 $\pm$ 0.072 0.930 $\pm$ 0.082 0.743 $\pm$ 0.112
    Qwen-2.5-0.5B 1.819 $\pm$ 0.027 45.091 $\pm$ 0.689 1.707 $\pm$ 0.045 0.572 $\pm$ 0.040 1.722 $\pm$ 0.033 2.492 $\pm$ 0.036 1.671 $\pm$ 0.064 1.492 $\pm$ 0.083 1.113 $\pm$ 0.114
Multimodal Models
    UniTime 0.405 $\pm$ 0.002 35.046 $\pm$ 0.176 0.361 $\pm$ 0.002 0.166 $\pm$ 0.001 0.319 $\pm$ 0.001 0.496 $\pm$ 0.005 0.314 $\pm$ 0.002 0.352 $\pm$ 0.002 0.409 $\pm$ 0.003
    Time-LLM (ETTh1) 0.458 $\pm$ 0.002 40.647 $\pm$ 0.155 0.440 $\pm$ 0.003 0.191 $\pm$ 0.000 0.371 $\pm$ 0.000 0.499 $\pm$ 0.002 0.399 $\pm$ 0.001 0.368 $\pm$ 0.001 0.424 $\pm$ 0.001
TS Foundation Models
    Lag-Llama 0.329 $\pm$ 0.004 29.074 $\pm$ 0.783 0.355 $\pm$ 0.007 0.181 $\pm$ 0.003 0.324 $\pm$ 0.003 0.272 $\pm$ 0.006 0.342 $\pm$ 0.006 0.386 $\pm$ 0.009 0.449 $\pm$ 0.012
    Chronos-Tiny 0.328 $\pm$ 0.001 26.277 $\pm$ 0.416 0.400 $\pm$ 0.002 0.144 $\pm$ 0.003 0.297 $\pm$ 0.002 0.229 $\pm$ 0.002 0.290 $\pm$ 0.005 0.382 $\pm$ 0.002 0.440 $\pm$ 0.003
    Chronos-Mini 0.341 $\pm$ 0.001 27.560 $\pm$ 0.394 0.412 $\pm$ 0.002 0.147 $\pm$ 0.002 0.302 $\pm$ 0.002 0.248 $\pm$ 0.002 0.305 $\pm$ 0.004 0.378 $\pm$ 0.003 0.436 $\pm$ 0.004
    Chronos-Small 0.328 $\pm$ 0.002 25.217 $\pm$ 0.349 0.388 $\pm$ 0.003 0.144 $\pm$ 0.002 0.287 $\pm$ 0.002 0.248 $\pm$ 0.003 0.290 $\pm$ 0.003 0.358 $\pm$ 0.005 0.412 $\pm$ 0.006
    Chronos-Base 0.672 $\pm$ 0.003 29.202 $\pm$ 0.353 0.702 $\pm$ 0.002 0.143 $\pm$ 0.002 1.023 $\pm$ 0.006 0.261 $\pm$ 0.003 0.299 $\pm$ 0.004 1.643 $\pm$ 0.009 2.187 $\pm$ 0.012
    Chronos-Large 0.326 $\pm$ 0.002 24.459 $\pm$ 0.380 0.385 $\pm$ 0.002 0.138 $\pm$ 0.002 0.288 $\pm$ 0.002 0.249 $\pm$ 0.002 0.295 $\pm$ 0.003 0.362 $\pm$ 0.003 0.417 $\pm$ 0.004
    TimeGEN 0.354 $\pm$ 0.000 34.432 $\pm$ 0.165 0.402 $\pm$ 0.000 0.176 $\pm$ 0.000 0.308 $\pm$ 0.000 0.279 $\pm$ 0.000 0.324 $\pm$ 0.000 0.377 $\pm$ 0.000 0.431 $\pm$ 0.000
    Moirai-small 0.565 $\pm$ 0.031 33.878 $\pm$ 0.421 0.429 $\pm$ 0.005 0.671 $\pm$ 0.146 0.468 $\pm$ 0.076 0.566 $\pm$ 0.017 1.204 $\pm$ 0.271 0.376 $\pm$ 0.009 0.426 $\pm$ 0.012
    Moirai-base 0.624 $\pm$ 0.013 33.138 $\pm$ 0.378 0.410 $\pm$ 0.006 0.600 $\pm$ 0.053 0.680 $\pm$ 0.028 0.690 $\pm$ 0.019 1.147 $\pm$ 0.099 0.375 $\pm$ 0.010 0.441 $\pm$ 0.013
    Moirai-large 0.520 $\pm$ 0.006 27.016 $\pm$ 0.843 0.414 $\pm$ 0.004 0.155 $\pm$ 0.004 0.260 $\pm$ 0.003 0.751 $\pm$ 0.015 0.276 $\pm$ 0.008 0.337 $\pm$ 0.007 0.397 $\pm$ 0.010
Statistical Models
    ARIMA 0.480 $\pm$ 0.006 25.581 $\pm$ 0.447 0.399 $\pm$ 0.006 0.160 $\pm$ 0.002 0.517 $\pm$ 0.012 0.522 $\pm$ 0.013 0.706 $\pm$ 0.026 0.354 $\pm$ 0.007 0.403 $\pm$ 0.010
    ETS 0.522 $\pm$ 0.009 31.849 $\pm$ 0.598 0.407 $\pm$ 0.009 0.228 $\pm$ 0.010 0.682 $\pm$ 0.018 0.571 $\pm$ 0.019 0.855 $\pm$ 0.035 0.453 $\pm$ 0.012 0.479 $\pm$ 0.015
    Exp-Smoothing 0.603 $\pm$ 0.013 34.050 $\pm$ 0.336 0.571 $\pm$ 0.021 0.334 $\pm$ 0.013 0.743 $\pm$ 0.018 0.557 $\pm$ 0.019 0.899 $\pm$ 0.035 0.673 $\pm$ 0.038 0.782 $\pm$ 0.053
C.4 Inference Time
Figure 13: Inference time in seconds, for all baselines, averaged over all tasks. Several quantitative methods are much faster on average than LLM-based methods. However, there are significant differences in inference time between the LLM-based forecasters: for the Llama models, llmp takes about an order of magnitude more time to run on average than direct prompt .
Figure 13 provides the inference time of all tested models on the benchmark. Note that these values have not been normalized based on the computing resources made available to each model during inference; please refer to Appendix D for information about how much compute resources were allocated to each of them.
C.5 Significant failures per model
@@
With Context
direct prompt
Llama-3.1-405B-Inst
0
Llama-3-70B-Inst
1
Llama-3-8B-Inst
2
Mixtral-8x7B-Inst
7
Qwen-2.5-7B-Inst
7
Qwen-2.5-1.5B-Inst
10
Qwen-2.5-0.5B-Inst
7
GPT-4o
5
GPT-4o-mini
0
@@
Llama-3-70B-Inst
18
Llama-3-70B
0
Llama-3-8B-Inst
12
Llama-3-8B
0
Mixtral-8x7B-Inst
1
Mixtral-8x7B
0
Qwen-2.5-7B-Inst
107
Qwen-2.5-7B
27
Qwen-2.5-1.5B-Inst
95
Qwen-2.5-1.5B
102
Qwen-2.5-0.5B-Inst
102
Qwen-2.5-0.5B
111